# An End-to-End Model for Photo-Sharing Multi-modal Dialogue Generation

### v1

@@ 4. Conclusion
In this paper, we proposed the first end-to-end photo-sharing multi-modal dialogue model, which consists of three modules: text dialogue generator, image perceptron and image generator. We exploited a large language model as the backbone of the text dialogue generator. For the image perceptron, Q-Former equipped the large language model with the ability to see images in the input end. For the image generator, we applied the stable diffusion model to enable the large language model to draw images in the output end. Particularly, we proposed the dynamic vocabulary transformation matrix and the straight-through gumbel-softmax technique to align the large language model and stable diffusion model and achieve end-to-end gradient propagation. Finally, we conducted experiments on PhotoChat and DialogCC datasets to verify the effectiveness of our proposed end-to-end model. Compared with two pipeline models, the end-to-end model gained state-of-the-art performances for photo-sharing multi-modal dialogue generation.
Acknowledgements.
This work was supported by the stability support plan for Shenzhen colleges and universities.

### v2

@@ V Conclusion
In this paper, we proposed the first end-to-end photo-sharing multi-modal dialogue model, which consists of three modules: text dialogue generator, image perceptron and image generator. We exploited an LLM as the backbone of the text dialogue generator. For the image perceptron, Llama-Vision equipped the LLM with the ability to see images in the input end. For the image generator, we applied the stable diffusion model to enable the LLM to draw images in the output end. Particularly, we proposed the dynamic vocabulary transformation matrix and the straight-through gumbel-softmax technique to align the LLM and stable diffusion model and achieve end-to-end gradient propagation. Finally, we conducted experiments on PhotoChat and DialogCC datasets to verify the effectiveness of our end-to-end model. Compared with pipeline models, the end-to-end model gained state-of-the-art performances for photo-sharing multi-modal dialogue generation.